OpenStudy (studygurl14):

calculus@michele_laino @satellite73

OpenStudy (studygurl14):

@Michele_Laino @satellite73

OpenStudy (anonymous):

\[\frac{d}{dx}f(2x)=2f'(2x)\]

OpenStudy (studygurl14):

how do you know?

OpenStudy (anonymous):

chain rule, it is always true

OpenStudy (anonymous):

like \[\frac{d}{dx}\sin(2x)=2\cos(2x)\]

OpenStudy (studygurl14):

okay.

OpenStudy (anonymous):

so you know \[2f'(2x)=f'(x)\] since they tell you they are equal

OpenStudy (studygurl14):

yes, ok

OpenStudy (studygurl14):

And if f'(1) = 1....then can you substitute that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you almost had it there

OpenStudy (studygurl14):

2f'(2x) = 1 f'(2x) = 1/2

OpenStudy (anonymous):

\[2f'(2\times 1)=f'(1)=1\]

OpenStudy (anonymous):

you should replace x by 1 in both spots

OpenStudy (studygurl14):

f'(2) = 1/2

OpenStudy (anonymous):

looks good to me

OpenStudy (studygurl14):

Wow, thanks! I didn't realize it would be so easy. Thanks!

OpenStudy (anonymous):

lol everything is easy when you know how

OpenStudy (anonymous):

btw can you think of a function \(f\) where \[f'(2x)=f'(x)\]?

OpenStudy (anonymous):

hint: is is a pretty common function, but not a polynomial, or rational function

OpenStudy (studygurl14):

hmmm....x^0

OpenStudy (studygurl14):

?

OpenStudy (anonymous):

\(x^0=1\)

OpenStudy (anonymous):

i suppose a constant function is one example, but i was looking for an example of a function that is not constant

OpenStudy (studygurl14):

1/x ?

OpenStudy (studygurl14):

no, nvm. you said it wasn't rational

OpenStudy (studygurl14):

idk, what is it?

OpenStudy (anonymous):

ask @freckles

OpenStudy (freckles):

\[u(2x)-u(x)=0 \\ u(x)=C \cdot n^{x } \\ u(2x)=C \cdot n^{2x} \\ \\ C n^{2x}-C n^{x}=0 \\ \text{ assume } C \neq 0, n \neq 0 \\ n^{2x}-n^{x}=0 \\ n^{x}(n^x-1)=0 \\ n^x=0 \text{ or } n^x=1 \\ \implies n=1 \\ \\ \text{ so } u(x)=C \\ f'(x)=C \\ \text{ since I replaced } f'(x) \text{ with } u(x) \\ f(x)=Cx+D \] even though I said C cannot be 0 C=0 does also work anyways this is a linear function and satellite asked for a non-polynomial and for some reason I cannot think that way lol I been playing with recurrence relations too much I could instead try to use a generating function to solve u(2x)-u(x)=0 and see if I get a different result

OpenStudy (anonymous):

did my picture help?

OpenStudy (freckles):

oh log function

OpenStudy (anonymous):

that one will work i am not sure any others will

OpenStudy (freckles):

I'm not seeing which log function for f satisfies f'(2x)=f'(x) :(

OpenStudy (freckles):

did you mean \[(f(2x))'=f'(x)\] @satellite73 \[2 f'(2x)=f'(x) \\ f'(x)=\frac{k}{x} \\ f''(2x)=\frac{k}{2x} \\ 2 \cdot \frac{k}{2x}=\frac{k}{x} \text{ is true } \\ f'(x)=\frac{k}{x} \implies f(x)= k \ln(x)+C\]

OpenStudy (freckles):

oops that '' ignore that should be a single mark

OpenStudy (anonymous):

in the link it says \[\frac{d}{dx}f(2x)=f'(x)\]

OpenStudy (freckles):

yes but I have been trying to solve f'(2x)=f'(x) forever now and get log somehow from it :p

OpenStudy (anonymous):

so since \[\log(2x)=\log(2)+\log(x)\] you have \[\log'(2x)=\log'(x)\]

OpenStudy (anonymous):

oh i see

OpenStudy (freckles):

you like to abuse things I see

OpenStudy (anonymous):

i do

OpenStudy (anonymous):

if i had clue one as to how to solve a differential equation, maybe i could prove that that is the only solution

OpenStudy (freckles):

\[2 f'(2x)=f'(x)\] i was thing about treating it similar to a recurrence relation and see what happens

OpenStudy (anonymous):

ooh i see

OpenStudy (anonymous):

so if \[2f(2x)=f(x)\] you have \[f(x)=\frac{k}{x}\]right?

OpenStudy (freckles):

yes

OpenStudy (anonymous):

so it is the log in every case

OpenStudy (anonymous):

learn something new every day

OpenStudy (anonymous):

i learned something else just now